# Symbolic computation of optimal systems of subalgebras of three- and four-dimensional real Lie algebras

### v1

@@ 2.1 One-dimensional Lie subalgebras
where $\mathcal{P}=\{f^{1},f^{2},t\}$, is a 2-family of one-dimensional Lie subalgebras which has to be considered indistinguishable from $X$. On the contrary, the family
$Y=f^{1}\Xi_{1}+f^{1}(f^{2})^{2}\Xi_{2}+f^{1}f^{2}\Xi_{3},$
even if the vector $(f^{1},f^{1}(f^{2})^{2},f^{1}f^{2},0)$ possesses three nonzero components, does not correspond to a $3$-family; in fact, we have $\mathcal{P}=\{f^{1},f^{2}\}$, and the rank of the Jacobian matrix of the coordinates of $Y$ with respect to the elements of $\mathcal{P}$ is 2.
@@ 2.2 Higher dimensional Lie subalgebras
According to Definition 2, we have $2^{r}-1$ possible one-dimensional $p$-families of Lie subalgebras, each identified by the tuple $X=(f^{1}d_{1},\ldots,f^{r}d_{r})$. Hence, the action of an inner automorphism on a $p$-family can be transferred to the corresponding tuple. Furthermore, although such $p$-families can be identified with the integers $c_{X}$, within the package they are sorted according to short-lexicographic (slex) ordering. The position of a $p$-family in this ordering is used to label the vertex representing it in the oriented multigraph $\mathcal{G}(\mathcal{L})$. Then, the vertices of such a graph are the one-dimensional $p$-families of Lie subalgebras of $\mathcal{L}$, and the edges are given by the action of the inner automorphisms on the vertices.
More technically, let $X=\{f^{1}d_{1},\ldots,f^{r}d_{r}\}$ and $Y=\{g^{1}d^{\prime}_{1},\ldots,g^{r}d^{\prime}_{r}\}$ be two families of one-dimensional Lie subalgefbras of $\mathcal{L}$. We already said that $X$ and $Y$ belong to the relation $\mathcal{R}$ if the action of some inner automorphism $A$ on each element of $X$ produces a $p$-family $Z$ of Lie subalgebras such that $c_{Z}=c_{Y}$. During the computation, if it is found that $X\,\mathcal{R}\,Y$, the adjacency matrix of $\mathcal{G}(\mathcal{L})$ is updated and it is checked whether $Y\,\mathcal{R}\,X$ too. If $X\,\mathcal{R}\,Y$ only, in the graph the vertices corresponding to $X$ and $Y$ are connected by a directed edge, whereas if it is also $Y\,\mathcal{R}\,X$ by an undirected edge. At the end of all these checks, the multigraph can be represented, and for each connected component a representative is selected: the vertex representing the smallest (in the slex ordering) $p$-family with maximal indegree. The set of these representatives yields the optimal system of families of one-dimensional Lie subalgebras.
@@ 6 Conclusions
The algorithmic construction and the analysis of the optimal systems in this paper have been performed using the package SymboLie running on Wolfram Mathematica™. The implementation of the algorithm relies on the definition of a $p$-family of Lie subalgebras; this implies that the closure of a vector subspace with respect to the Lie bracket can be ascertained without solving quadratic equations. Although this approach may leave out some Lie subalgebras [15], in most of the cases this strategy is effective and provides the complete results. Then, introducing a relation between families of subalgebras induced by the inner automorphisms (this relation is in general a preorder), we partition the families of Lie subalgebras and represent them by a graph: the simplest representatives of the connected components of this graph give the list of optimal systems of families of Lie subalgebras. The results presented in this paper, dealing with all three- and four-dimensional Lie algebras, represent a good test for the program. The analysis of higher dimensional Lie algebras is currently in progress, as well as the implementation of some strategies to speed up the algorithm. Anyway, we are confident that such a package in its present version can be useful to researchers.

### v2

@@ 2.1 One-dimensional Lie subalgebras
where $\mathcal{P}=\{f^{1},f^{2},t\}$, is a 2-family of one-dimensional Lie subalgebras which has to be considered indistinguishable from $X$. On the contrary, the element
$Y=f^{1}\Xi_{1}+f^{2}\Xi_{2}+f^{3}\Xi_{3},$
where the components $f^{1}$, $f^{2}$ and $f^{3}$ satisfy the relation $f^{1}f^{2}=(f^{3})^{2}$, does not correspond to a 3-family. Indeed, although the vector $(f^{1},f^{2},f^{3},0)$ possesses three nonzero components, we do not have a $3$-family because the three components are not independent.
@@ 2.2 Higher dimensional Lie subalgebras
According to Definition 2, we have $2^{r}-1$ possible one-dimensional $p$-families of Lie subalgebras, each identified by the tuple $X=(f^{1}s_{1},\ldots,f^{r}s_{r})$. Hence, the action of an inner automorphism on a $p$-family can be transferred to the corresponding tuple. Furthermore, although such $p$-families can be identified with the integers $c_{X}$, within the package they are sorted according to short-lexicographic (slex) ordering. The position of a $p$-family in this ordering is used to label the vertex representing it in the oriented multigraph $\mathcal{G}(\mathcal{L})$. Then, the vertices of such a graph are the one-dimensional $p$-families of Lie subalgebras of $\mathcal{L}$, and the edges are given by the action of the inner automorphisms on the vertices.
More technically, let $X=\{f^{1}s_{1},\ldots,f^{r}s_{r}\}$ and $Y=\{g^{1}s^{\prime}_{1},\ldots,g^{r}s^{\prime}_{r}\}$ be two families of one-dimensional Lie subalgebras of $\mathcal{L}$. We already said that $X$ and $Y$ belong to the relation $\mathcal{R}$ if the action of some inner automorphism $A$ on each element of $X$ produces a $p$-family $Z$ of Lie subalgebras such that $c_{Z}=c_{Y}$. During the computation, if it is found that $X\,\mathcal{R}\,Y$, the adjacency matrix of $\mathcal{G}(\mathcal{L})$ is updated and it is checked whether $Y\,\mathcal{R}\,X$ too. If $X\,\mathcal{R}\,Y$ only, in the graph the vertices corresponding to $X$ and $Y$ are connected by a directed edge, whereas if it is also $Y\,\mathcal{R}\,X$ by an undirected edge. At the end of all these checks, the multigraph can be represented, and for each connected component a representative is selected: the vertex representing the smallest (in the slex ordering) $p$-family with maximal indegree. The set of these representatives yields the optimal system of families of one-dimensional Lie subalgebras.
@@ 6 Conclusions
In this paper, the algorithmic construction and the analysis of the optimal systems have been performed using the package SymboLie running on Wolfram Mathematica™. The implementation of the algorithm relies on the definition of $p$-family of Lie subalgebras; this implies that the closure of a vector subspace with respect to the Lie bracket can be ascertained without solving quadratic equations. We are aware that this approach may leave out some Lie subalgebras [15], i.e., those Lie subalgebras not representing $p$-families. Anyway, in most of the cases this strategy is effective and provides the complete results. Then, introducing a relation between families of subalgebras induced by the inner automorphisms (this relation is in general a preorder), we partition the families of Lie subalgebras and represent them by a graph: the simplest representatives of the connected components of this graph give the list of optimal systems of families of Lie subalgebras. The results presented in this paper, dealing with all three- and four-dimensional Lie algebras, deeply investigated in [23], represent a good test for the program. The results confirm those provided in [23], with two exceptions. For algebra $A_{3,8}$, in [23] the one-dimensional optimal system contains the subalgebra $\{\Xi_{1}+\Xi_{3}\}$ which is not a family according to our definition, whereas SymboLie produces the family $\{\Xi_{1}+\alpha_{1}\Xi_{3}\}$, $\alpha_{1}=\pm 1$. In the case of the algebra $A_{4,5}^{a,b}$, the results given in [23] miss the family $\{\Xi_{1}+\alpha_{1}\Xi_{2}\}$, that is listed by SymboLie. It is easy to recognize that this family must belong to the optimal system because it is invariant with respect to all the inner automorphisms of the Lie algebra.
The analysis of higher dimensional Lie algebras is currently in progress, as well as the implementation of some strategies to speed up the algorithm. Anyway, we are confident that such a package in its present version can be useful to researchers.